# VITA: Variational Pretraining of Transformers for Climate-Robust Crop Yield Forecasting

### v2

@@ Abstract
Accurate crop yield forecasting is essential for global food security. However, current AI models systematically underperform when yields deviate from historical trends. We attribute this to the lack of rich, physically grounded datasets directly linking atmospheric states to yields. To address this, we introduce VITA (Variational Inference Transformer for Asymmetric data), a variational pretraining framework that learns representations from large satellite-based weather datasets and transfers to the ground-based limited measurements available for yield prediction. VITA is trained using detailed meteorological variables as proxy targets during pretraining and learns to predict latent atmospheric states under a seasonality-aware sinusoidal prior. This allows the model to be fine-tuned using limited weather statistics during deployment. Applied to 763 counties in the U.S. Corn Belt, VITA achieves state-of-the-art performance in predicting corn and soybean yields across all evaluation scenarios, particularly during extreme years, with statistically significant improvements (paired t-test, $p<0.0001$). Importantly, VITA outperforms prior frameworks like GNN-RNN without soil data, and bigger foundational models (e.g., Chronos-Bolt) with less compute, making it practical for real-world use—especially in data-scarce regions. This work highlights how domain-aware AI design can overcome data limitations and support resilient agricultural forecasting in a changing climate.
@@ Introduction
Climate change is transforming agriculture, with extreme weather causing billions in annual crop losses (Lobell, Schlenker, and Costa-Roberts 2011). In 2012, U.S. drought reduced corn yields by 13%, while 2019 flooding prevented planting on 19.4 million acres (USDA National Agricultural Statistics Service 2013; USDA Farm Service Agency 2019). Accurate yield prediction under such volatility is critical for agricultural risk management and long-term food security (Beddington 2010). Yet current operational models—including regression-based approaches from USDA ERS (Westcott and Jewison 2013) and USDA NASS (NASS 2023)—often fail when yields diverge from historical trends.
@@
Third, general-purpose time-series pretraining methods like SimMTM (Dong et al. 2023), PatchTST (Wu et al. 2023), and foundational models, such as, Chronos (Ansari et al. 2024) assume consistent input features between pretraining and fine-tuning. This is suboptimal weather domains, where pretraining can leverage rich satellite datasets with dozens of variables (e.g., 31 meteorological variables from NASA POWER (NASA 2024)), but fine-tuning must rely on smaller, accessible subsets (e.g., 6 basic weather variables from Khaki, Wang, and Archontoulis (2019)). We term this the data asymmetry problem, in which pretraining and operational feature sets fundamentally differ—an issue largely unaddressed by existing AI approaches.
@@
The remainder of this paper is organized as follows: Related Work reviews prior methods; Methodology details our variational framework, Experiments presents our experimental setup, Results analyzes performance across various conditions, and Discussion discusses implications and limitations.
@@
Researchers have proposed various different approaches to for crop yield forecasting, including mechanistic modeling, CNN-RNN architectures, graph neural networks, Deep Gaussian Processes, and Vision Transformers (Keating et al. 2003; Khaki, Wang, and Archontoulis 2019; Fan et al. 2021; You et al. 2017; Lin et al. 2024). Multi-modal approaches integrate satellite imagery, weather data, soil surveys, management records, and vegetation indices (Wu et al. 2021; Sun et al. 2019; Gandhi, Petkar, and Armstrong 2016; Oliveira et al. 2018; Basir et al. 2021; Hasan, Roozbehani, and Dahleh 2024; Ferraz et al. 2024; Cao, Ma, and Zhang 2022). However, extensive data requirements limit deployment in regions with limited agricultural monitoring infrastructure, and many approaches rely on temporally short regional datasets that may not capture sufficient weather variability for extreme events (Gandhi, Petkar, and Armstrong 2016; Lin et al. 2024; McFarland et al. 2020; Chu and Yu 2020).
@@
Recent work develops unsupervised representation learning through contrastive methods, transformer frameworks, and masked reconstruction (Franceschi, Dieuleveut, and Jaggi 2019; Yue et al. 2022; Zerveas et al. 2021; Dong et al. 2023; Wu et al. 2023; Woo et al. 2022). These methods achieve strong forecasting performance but do not explicitly model data asymmetry, which is critical for weather based yield forecasting.
@@
Variational autoencoders and have been applied to weather prediction, climate forecasting, and agricultural data generation (Kingma and Welling 2013; Higgins et al. 2017; Kwok and Qi 2021; Wang et al. 2024; Palma et al. 2025; Razavi et al. 2024). However, existing variational methods are not used for asymmetric data learning. Our work is applicable for this particular problem and learns latent state without input reconstruction and with a seasonality-aware sinusoidal prior.
@@
Figure 2: Graphical model depicting data structure of pretraining and prediction phases in VITA.
@@
We consider two datasets: (1) a pretraining dataset $\mathcal{D}_{w}=\{(x_{i},z_{i})\}_{i=1}^{N_{w}}$, where both basic and detailed weather states are observed for each 364-week non-overlapping window $i$ across the NASA POWER grid, but no yield information is available; and (2) a finetuning dataset $\mathcal{D}_{y}=\{(x_{j},y_{j},y_{\text{past},j})\}_{j=1}^{N_{y}}$, where detailed states $z_{j}$ remain latent for each 364-week overlapping sequence $j$ across U.S. Corn Belt counties. Each training example uses weather from years [t6,t][t-6,t] and historical yields from [t6,t1][t-6,t-1] to predict yield $y_{c,t}$ for county $c$. Test-year yields are strictly held out during training and additional details are provided in the Appendix. We model each county independently as a 1D temporal sequences and spatial context enters only via (latitude,longitude) features concatenated to each timestep (Equation 1).
@@
$\hat{y}_{t}$ $=\mathrm{MLP}_{y}\!\big([\,z_{\text{agg}},\,y_{\text{past},t}\,]\big)$ (6)
@@
We pretrain on NASA POWER dataset (1984-2022) using progressive feature-wise masking, starting with $k=10$ masked features and increasing by 1 every 2 epochs until 25 out of 31 features are masked. The pretraining objective balances reconstruction with regularization:
@@
The first term maximizes the Gaussian likelihood, which encourages the posterior distribution $q_{\phi}(z_{i}\mid x_{i})$ to accurately predict the observed detailed weather state $z_{i}$. The second term is the regularizer preventing overfitting by imposing a prior structure.
We investigate two prior distributions: standard normal $p_{\theta}(z)\sim\mathcal{N}(0,I)$ and sinusoidal prior $p_{\theta}(z)\sim\mathcal{N}(\mu_{\sin}=A\sin(\theta\cdot\text{pos}+\theta_{0}),\sigma^{2}I)$ to capture seasonal patterns. Sinusoidal prior has additional prior parameters that are learned during both pretraining and finetuning. This prior explicitly models the periodicity in weather variables, allowing a more structured latent space.
@@
In Appendix, we empirically validated this deterministic relationship, training a small MLP to reconstruct basic weather variables from detailed ones with near-perfect accuracy ($R^{2}>0.9999$). This enables us to model $p(x_{j}|z_{j})\approx 1$ and derive the simplified variational objective:
@@
We compare against several types of baselines. (1) Non-deep learning methods: OLS linear regression with agronomically-motivated features following USDA Economic Research Service (ERS) methodology (Westcott and Jewison 2013), and XGBoost (Chen and Guestrin 2016). (2) Deep learning methods: CNN-RNN (Khaki, Wang, and Archontoulis 2019) and GNN-RNN (Fan et al. 2021). (3) Masked time series pretraining: SimMTM (Dong et al. 2023). (4) Pre-trained foundational time series models: Chronos-Bolt-tiny-9M (Ansari et al. 2024) with full fine-tuning. To isolate the effect of variational pretraining from architecture, we also include T-BERT, which is identical to VITA but trained with a standard MSE reconstruction loss. Notably, XGBoost, GNN-RNN, and CNN-RNN have access to soil data, while the pretrained transformer models and OLS do not. All models use identical 364 week windows and train-test splits.
@@
Figure 3: Mean crop yield in bushels / acre (bu/ac) 763 US Corn Belt counties showing extreme weather years as sharp deviations from historical patterns.
@@
We pretrain on the NASA POWER dataset (NASA 2024), comprising 39 years (1984–2022) of climate data at 0.5° resolution across 116 grids over the Americas. It includes 31 meteorological variables aggregated weekly (100K sequences). However, its coarse resolution is unsuitable for county-level yield forecasting, as the median U.S. county spans 622 square miles (U.S. Census Bureau 2025). We therefore evaluate on the dataset of Khaki, Wang, and Archontoulis (2019), containing weather, soil, and county-level corn and soybean yields for 763 Corn Belt counties (1982–2018) (U.S. Department of Agriculture, National Agricultural Statistics Service 2023). Corn (C4) and soybean (C3) represent distinct physiological and weather-sensitivity regimes and jointly account for over 60% of U.S. row-crop acreage (Williams and Pounds-Barnett 2024).
The Khaki, Wang, and Archontoulis (2019) weather data include six ground-based weekly weather measurements: (1) minimum temperature, (2) maximum temperature, (3) solar radiation, (4) precipitation, (5) snow water equivalent, and (6) vapor pressure, and 11 soil properties averaged over county areas. We exclude soil data to test deployment in data-sparse regions. This presents a key challenge as pretraining contains satellite measurement of 31 meteorological variables, while fine-tuning uses only 6 ground-based variables averaged over county areas.
@@
We also conduct an early-season forecasting experiment using our top four models (two VITA variants, T-BERT, and SimMTM) and OLS, truncating weather data at the end of July in the final year (week 30). This design mirrors operational forecasting systems such as USDA NASS (NASS 2023), which issue yield projections prior to harvest. Our OLS baseline follows the USDA Economic Research Service (ERS) framework (Westcott and Jewison 2013), which is conceptually similar to NASS models that also rely on regression-driven estimates grounded in observed weather and crop conditions. The OLS baseline normally incorporates July–August temperature and precipitation for soybean (Appendix OLS Baseline: Domain-Engineered Features), but for consistency across methods we limit it to July in this experiment.
@@ Results
Figure 4: Individual R² comparisons on extreme years showing consistent improvements of VITA-Sinusoidal over other baselines.
@@
Table 3: Early-season forecasting on extreme years. Full: 7 years weather through end of season (52 weeks). Week 30: cutoff at end of July (30 weeks of final year).
@@
Operational viability is confirmed through two stress tests. Weather cut off at July (week 30 of final year) VITA-Sinusoidal maintains strong performance, achieving 0.689 R² on corn and 0.560 R² – nearly 3.0× and 1.5× better than OLS baselines, respectively (Table 3). The performance drop for soybean is particularly sharp across all models, as its late pod-filling stage makes yields highly sensitive to August rainfall and temperature stress. (Westcott and Jewison 2013)
@@
Standard years (2014-2018) validate that extreme-weather optimization doesn’t compromise normal conditions. VITA achieves 0.827 R² for corn and 0.833 R² for soybean with 15-year training—improvements of +4.6% and +0.2% over T-BERT respectively. The contrast with extreme years (+10.5% corn improvement) supports our core hypothesis: variational uncertainty modeling provides greatest value precisely when predictions are hardest. With 30-year training, corn performance rises to 0.837 R² (+7.3% over T-BERT), suggesting additional historical data helps but yields diminishing returns compared to better representations.
@@
Table 5: VITA pretraining ablation across 27 hyperparameter (random seed 1234).
@@ Discussion
Current USDA operational forecasts—including ERS regression models (Westcott and Jewison 2013) and NASS survey-driven approaches (NASS 2023)—rely on simpler models with hand-crafted features, achieving 0.227 R² for extreme corn years with our OLS baseline. VITA’s 3.2× improvement (0.729 R²) translates to substantial impact: 11.4 bu/ac RMSE reduction over OLS and 2.0 bu/ac over T-BERT across 88.7M Corn Belt acres (National Corn Growers Association 2024). At \$4.70/bushel (DTN Progressive Farmer 2025), this translates to \$4.75B and \$800M in value, respectively. This accuracy is critical for the Federal Crop Insurance Program managing billions in premiums (USDA Economic Research Service 2024) and for policy responses during droughts. Deployment requires minimal infrastructure (single GPU, 2.5 hours training) with only public data, enabling integration into existing agricultural statistics systems.
@@
The decoder-free variational framework generalizes beyond agriculture to any setting with rich sensors at training and sparse sensors at inference (e.g., ICU vitals vs. labs, IoT vs. industrial telemetry). Current evaluation focuses on U.S. Corn Belt corn and soybean (Khaki, Wang, and Archontoulis 2019; Fan et al. 2021); extending to other crops and regions remains future work. We commit to open-sourcing implementation, preprocessing scripts, model weights, and documentation for new region adaptation. Data privacy concerns are negligible as we use aggregated public meteorological data, ensuring equitable access through transparency.

### v3

@@ Abstract
Accurate crop yield forecasting is essential for global food security. However, current AI models systematically underperform when yields deviate from historical trends. We attribute this to the lack of rich, physically grounded datasets directly linking atmospheric states to yields. To address this, we introduce VITA (Variational Inference Transformer for Asymmetric data), a variational pretraining framework that learns representations from large satellite-based weather datasets and transfers to the ground-based limited measurements available for yield prediction. VITA is trained using detailed meteorological variables as proxy targets during pretraining and learns to predict latent atmospheric states under a seasonality-aware sinusoidal prior. This allows the model to be fine-tuned using limited weather statistics during deployment. Applied to 763 counties in the U.S. Corn Belt, VITA achieves state-of-the-art performance in predicting corn and soybean yields across all evaluation scenarios, particularly during extreme years, with statistically significant improvements (paired t-test, $p<0.0001$). Importantly, VITA outperforms prior frameworks like GNN-RNN without soil data, and larger foundational models (e.g., Chronos-Bolt) with less compute, making it practical for real-world use—especially in data-scarce regions. This work highlights how domain-aware AI design can overcome data limitations and support resilient agricultural forecasting in a changing climate.
Codehttps://github.com/neehan/VITA
@@ Introduction
Climate change is transforming agriculture, with extreme weather causing billions in annual crop losses (Lobell, Schlenker, and Costa-Roberts 2011). In 2012, U.S. drought reduced corn yields by 13%, while 2019 flooding prevented planting on 19.4 million acres (USDA National Agricultural Statistics Service 2013; USDA Farm Service Agency 2019). Accurate yield prediction under such volatility is critical for agricultural risk management and long-term food security (Beddington 2010). Yet current operational models—including regression-based approaches from USDA ERS (Westcott and Jewison 2013)—often fail when yields diverge from historical trends.
@@
Third, general-purpose time-series pretraining methods like SimMTM (Dong et al. 2023), PatchTST (Wu et al. 2023), and foundational models, such as Chronos (Ansari et al. 2024) assume consistent input features between pretraining and fine-tuning. This is suboptimal in weather domains, where pretraining can leverage rich satellite datasets with dozens of variables (e.g., 31 meteorological variables from NASA POWER (NASA 2024)), but fine-tuning must rely on smaller, accessible subsets (e.g., 6 basic weather variables from Khaki, Wang, and Archontoulis (2019)). We term this the data asymmetry problem, in which pretraining and operational feature sets fundamentally differ—an issue largely unaddressed by existing AI approaches.
@@
The remainder of this paper is organized as follows: Related Work reviews prior methods; Methodology details our variational framework; Experiments presents our experimental setup; Results analyzes performance across various conditions; and Discussion examines implications and limitations.
@@
Researchers have proposed various different approaches for crop yield forecasting, including mechanistic modeling, CNN-RNN architectures, graph neural networks, Deep Gaussian Processes, and Vision Transformers (Keating et al. 2003; Khaki, Wang, and Archontoulis 2019; Fan et al. 2021; You et al. 2017; Lin et al. 2024). Multi-modal approaches integrate satellite imagery, weather data, soil surveys, management records, and vegetation indices (Wu et al. 2021; Sun et al. 2019; Gandhi, Petkar, and Armstrong 2016; Oliveira et al. 2018; Basir et al. 2021; Hasan, Roozbehani, and Dahleh 2024; Ferraz et al. 2024; Cao, Ma, and Zhang 2022). However, extensive data requirements limit deployment in regions with limited agricultural monitoring infrastructure, and many approaches rely on temporally short regional datasets that may not capture sufficient weather variability for extreme events (Gandhi, Petkar, and Armstrong 2016; Lin et al. 2024; McFarland et al. 2020; Chu and Yu 2020).
@@
Recent work develops unsupervised representation learning through contrastive methods, transformer frameworks, and masked reconstruction (Franceschi, Dieuleveut, and Jaggi 2019; Yue et al. 2022; Zerveas et al. 2021; Dong et al. 2023; Wu et al. 2023; Woo et al. 2022). These methods achieve strong forecasting performance but do not explicitly model data asymmetry, which is critical for weather-based yield forecasting.
@@
Variational autoencoders have been applied to weather prediction, climate forecasting, and agricultural data generation (Kingma and Welling 2013; Higgins et al. 2017; Kwok and Qi 2021; Wang et al. 2024; Palma et al. 2025; Razavi et al. 2024). However, existing variational methods are not used for asymmetric data learning. Our approach is tailored to this problem, learning latent states without input reconstruction and using a seasonality-aware sinusoidal prior.
@@
Figure 2: Graphical model showing the data structure of pretraining and prediction phases in VITA.
@@
We consider two datasets: (1) a pretraining dataset $\mathcal{D}_{w}=\{(x_{i},z_{i})\}_{i=1}^{N_{w}}$, where both basic and detailed weather states are observed for each 364-week non-overlapping window $i$ across the NASA POWER grid, but no yield information is available; and (2) a finetuning dataset $\mathcal{D}_{y}=\{(x_{j},y_{j},y_{\text{past},j})\}_{j=1}^{N_{y}}$, where detailed states $z_{j}$ remain latent for each 364-week overlapping sequence $j$ across U.S. Corn Belt counties. Each training example uses weather from years [t6,t][t-6,t] and historical yields from [t6,t1][t-6,t-1] to predict yield $y_{c,t}$ for county $c$. Test-year yields are strictly held out during training and additional details are provided in the Appendix. We model each county independently as a 1D temporal sequence and spatial context enters only via (latitude,longitude) features concatenated to each timestep (Equation 1).
@@
$\hat{y}_{t}$ $=\mathrm{MLP}_{y}\!\big([\,z_{\text{agg}},\,y_{\text{past}}\,]\big)$ (6)
@@
We pretrain on the NASA POWER dataset (1984-2022) using progressive feature-wise masking, starting with $k=10$ masked features and increasing by 1 every 2 epochs until 25 out of 31 features are masked. The pretraining objective balances reconstruction with regularization:
@@
The first term maximizes the Gaussian likelihood, which encourages the posterior distribution $q_{\phi}(z_{i}\mid x_{i})$ to accurately predict the observed detailed weather state $z_{i}$. The second term is a regularizer preventing overfitting by imposing a prior structure.
We investigate two prior distributions: standard normal $p_{\theta}(z)\sim\mathcal{N}(0,I)$ and sinusoidal prior $p_{\theta}(z)\sim\mathcal{N}(\mu_{\sin}=A\sin(\theta\cdot\text{pos}+\theta_{0}),\sigma^{2}I)$ to capture seasonal patterns. The Sinusoidal prior has additional prior parameters that are learned during both pretraining and finetuning. This prior explicitly models the periodicity in weather variables, allowing a more structured latent space.
@@
In the Appendix, we empirically validated this deterministic relationship, training a small MLP to reconstruct basic weather variables from detailed ones with near-perfect accuracy ($R^{2}>0.9999$). This enables us to model $p(x_{j}|z_{j})\approx 1$ and derive the simplified variational objective:
@@
We compare against several types of baselines. (1) Non-deep learning methods: OLS linear regression with agronomically-motivated features following USDA Economic Research Service (ERS) methodology (Westcott and Jewison 2013), and XGBoost (Chen and Guestrin 2016). (2) Deep learning methods: CNN-RNN (Khaki, Wang, and Archontoulis 2019) and GNN-RNN (Fan et al. 2021). (3) Masked time series pretraining: SimMTM (Dong et al. 2023). (4) Pre-trained foundational time series models: Chronos-Bolt-tiny-9M (Ansari et al. 2024) with full fine-tuning. To isolate the effect of variational pretraining from architecture, we also include T-BERT, which is identical to VITA but trained with a standard MSE reconstruction loss. Notably, XGBoost, GNN-RNN, and CNN-RNN have access to soil data, while the pretrained transformer models and OLS do not. All models except the OLS use identical temporal windows and train-test splits.
@@
Figure 3: Mean crop yield in bushels per acre (bu/ac) 763 US Corn Belt counties showing extreme weather years as sharp deviations from historical patterns.
@@
We pretrain on the NASA POWER dataset (NASA 2024), comprising 39 years (1984–2022) of climate data at 0.5° resolution across 116 grids over the Americas. It includes 31 meteorological variables aggregated weekly (100K sequences). However, its resolution is too coarse for county-level yield forecasting, as the median U.S. county spans 622 square miles (U.S. Census Bureau 2025). We therefore evaluate on the dataset of Khaki, Wang, and Archontoulis (2019), containing weather, soil, and county-level corn and soybean yields for 763 Corn Belt counties (1982–2018) (U.S. Department of Agriculture, National Agricultural Statistics Service 2023). Corn (C4) and soybean (C3) represent distinct physiological and weather-sensitivity regimes and jointly account for over 60% of U.S. row-crop acreage (Williams and Pounds-Barnett 2024).
The Khaki, Wang, and Archontoulis (2019) weather data include six ground-based weekly weather measurements: (1) minimum temperature, (2) maximum temperature, (3) solar radiation, (4) precipitation, (5) snow water equivalent, and (6) vapor pressure, and 11 soil properties averaged over county areas. We exclude soil data to test deployment in data-sparse regions. This presents a key challenge, as pretraining contains satellite measurements of 31 meteorological variables, while fine-tuning uses only 6 ground-based variables averaged over county areas.
@@
We also conduct an early-season forecasting experiment using our top four models (two VITA variants, T-BERT, and SimMTM) and OLS, truncating weather data at the end of July in the final year (week 30). This design follows the USDA Economic Research Service (ERS) framework (Westcott and Jewison 2013), which is conceptually similar to NASS models that also rely on regression-driven estimates grounded in observed weather and crop conditions. The OLS baseline normally incorporates July–August temperature and precipitation for soybean (Appendix OLS Baseline: Domain-Engineered Features), but for consistency across methods we limit it to July in this experiment.
@@ Results
Figure 4: VITA-Sinusoidal shows consistent improvement over other baselines.
@@
Table 3: Early-season forecasting on extreme years for the top 4 models and OLS. Full: 7 years weather through end of season (52 weeks). Week 30: cutoff at end of July, ie. week 30 of the final year.
@@
Operational viability is confirmed through two stress tests. With weather cut off at the end of July (week 30 of the final year), VITA-Sinusoidal maintains strong performance, achieving 0.689 R² on corn and 0.560 R² – nearly 3.0× and 1.5× better than OLS baselines, respectively (Table 3). The performance drop for soybean is particularly sharp across all models, as its late pod-filling stage makes yields highly sensitive to August rainfall and temperature stress. (Westcott and Jewison 2013)
@@
Standard years (2014-2018) validate that extreme-weather optimization doesn’t compromise performance under normal conditions. VITA achieves 0.827 R² for corn and 0.833 R² for soybean with 15-year training—improvements of +4.6% and +0.2% over T-BERT respectively. The contrast with extreme years (+10.5% corn improvement) supports our core hypothesis: variational uncertainty modeling provides greatest value precisely when predictions are hardest. With 30-year training, corn performance rises to 0.837 R² (+7.3% over T-BERT), suggesting additional historical data helps but yields diminishing returns compared to better representations.
@@
Table 5: VITA pretraining ablation across 27 hyperparameter configurations (random seed 1234).
@@ Discussion
Current USDA operational forecasts—including ERS regression models (Westcott and Jewison 2013)—rely on simpler models with hand-crafted features, achieving 0.227 R² for extreme corn years with our OLS baseline. VITA’s 3.2× improvement (0.729 R²) translates to substantial impact: 11.4 bu/ac RMSE reduction over OLS and 2.0 bu/ac over T-BERT across 88.7M Corn Belt acres (National Corn Growers Association 2024). At \$4.70/bushel (DTN Progressive Farmer 2025), this translates to \$4.75B and \$800M in value, respectively. This accuracy is critical for the Federal Crop Insurance Program managing billions in premiums (USDA Economic Research Service 2024) and for policy responses during droughts. Deployment requires minimal infrastructure (single GPU, 2.5 hours training) with only public data, enabling integration into existing agricultural statistics systems.
@@
The decoder-free variational framework can ba applied beyond agriculture to any setting with rich sensors at training and sparse sensors at inference (e.g., ICU vitals vs. labs, IoT vs. industrial telemetry). Current evaluation focuses on U.S. Corn Belt corn and soybean (Khaki, Wang, and Archontoulis 2019; Fan et al. 2021); extending to other crops and regions remains future work. We open-sourced implementation, preprocessing scripts, model weights, and documentation for new region adaptation. Data privacy concerns are negligible as we use aggregated public meteorological data, ensuring equitable access through transparency.